OpenStudy (anonymous):

answer check: divide 24x^8 - 18x^6 + 30 / 6x = 4x^7 - 3x^5 + 5 ans.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I had it written correctly on paper forgot and left it out by mistake thnks

OpenStudy (anonymous):

ok just checking on answers now plse: subtract (2a^3 + 3b^3) - (6a^3 + 4a^2b - 4ab^2 + 3b^3 = -4a^3 - 4a^2b + 4ab^2 ans.

OpenStudy (anonymous):

ok just checking on answers now plse: subtract (2a^3 + 3b^3) - (6a^3 + 4a^2b - 4ab^2 + 3b^3 = -4a^3 - 4a^2b + 4ab^2 ans.

OpenStudy (anonymous):

\[\frac{24 x^8-18 x^6+30}{6 x}=4 x^7-3 x^5+\frac{5}{x} \]\[\left(4 x^7-3 x^5+\frac{5}{x}\right)6x=24 x^8-18 x^6+30 \]